North Carolina's Native-American population is 22,000 greater than New York's.New York's Native-American population is 187,000 less than Oklahoma's.If the total population of all three is 437,000,find each state's Native-American population.
Answer by mananth(16946)

North Carolina's Native-American population is 22,000 greater than New York's.
New York's Native-American population is 187,000 less than Oklahoma's.
total population of all three is 437,000,
..
Let Oklahoma's population be x
NY population is x-187000
North carolina's population is 22000 +x-187000
..
total = 437000
...
x+x-187000+22000+x-187000=437000
3x=187000+187000-22000+437000
3x=789000
/3
x= 263,000 Oklahoma.
New York 76,000
North Carolina 98,000
...
m.ananth@hotmail.ca